OpenStudy (anonymous):

On Sunday, a local hamburger shop sold a combined total of 258 hamburgers and cheeseburgers. The number of cheeseburgers sold was two times the number of hamburgers sold. How many hamburgers were sold on Sunday?

OpenStudy (anonymous):

Equations: Quantity Eq. : h + c = 448 Quantity Eq. : c = 3h

OpenStudy (anonymous):

Substitute to solve for "h": h + 3h = 448

OpenStudy (anonymous):

4h = 448 h = 112 (# of hamburgers) ---------- Substitute to solve for "c": 112 + c = 448 c = 336 (# of cheesburgers)

OpenStudy (anonymous):

@shifuyanli It says that it is incorrect

OpenStudy (anonymous):

but for future reference post this in the math section (:

OpenStudy (anonymous):

hmmm...what are your choices @Deldrickg ?

OpenStudy (anonymous):

There aren't any choices its a response question

OpenStudy (anonymous):

#112 of hamburgers

OpenStudy (anonymous):

112 of hamburgers is what it was asking sorry it is 112 I'm so stupid right now!!

OpenStudy (anonymous):

@Deldrickg

OpenStudy (anonymous):

it is 112